### v5

# Are Biological Systems More Intelligent Than Artificial Intelligence?

Abstract
Are biological self-organising systems more ‘intelligent’ than artificial intelligence (AI)? If so, why? I explore this through a mathematical lens which frames intelligence in terms of adaptability. I model systems as stacks of abstraction layers (Stack Theory) and compare them by how they delegate agentic control down their stacks, illustrating with examples of computational, biological, human military, governmental and economic systems. Contemporary AI rests on a static, human-engineered stack in which lower layers are static during deployment. Put provocatively, static stacks resemble inflexible bureaucracies, adapting only top-down. Biological stacks are more ‘intelligent’ because they delegate adaptation. Formally, I prove a theorem (The Law of the Stack) showing adaptability in higher layers requires sufficient adaptability in lower layers. Generalising bio-electric explanations of cancer as isolation from collective informational structures, I explore how cancer-like failures occur in non-biological systems when delegation is inadequate. This helps explain how to build more robust systems, by delegating control like the military doctrine of mission command. It also provides a design perspective on hybrid agents (e.g. organoids, systems involving both humans and AI): hybrid creation is a boundary-condition design problem in which human-imposed constraints prune low-level policy spaces to yield desired collective behaviour while preserving collective identity.
1 Introduction
Biological systems self-organise; their parts and the environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014]. They pursue homeostatic and reproductive goals. Some proactively learn [kukushkin2024], storing information about the past in order to anticipate future needs. To make accurate predictions in complex environments, a system must learn cause and effect [pearl2018, merker2005, barron2016]. I consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Throughout, “intelligence” is used in an efficiency-of-adaptation sense (sample and energy efficiency). For a comparison with other common definitions, see Electronic Supplementary Materials (ESM) Note S1.. In this view intelligence is a long-term adaptation hard-wired into the organism, that enables short-term adaptation during its lifetime [bennett2025thesis].
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [russell2005ai, bishop2006, sutton2018, goertzel2014]. Contemporary approaches to machine learning are often inefficient [bennettmaruyama2022a]. They can require vast quantities of data and energy to learn tasks a human can pick up after one or two examples. Even organoids [^2] By organoid I mean a collective of cells grown in a lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology. appear more ‘intelligent’ in this sense [khajehnejad2024]. For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency. That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions. This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [nakagaki2000]. These sorts of results are highly dependent upon the human imposed boundary conditions, but they raise an interesting question. Are biological self-organising systems more intelligent than artificial intelligence? If so, why?
Some say the key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [parr2024, Friston2010, friston2013]. It is true that active inference can improve efficiency [ororbia2022, ororbia2023, ororbia2023b]. Active inference has been applied in the context of the ARC-AGI benchmark [chollet2019], which to some extent tests sample efficiency. Active inference systems performed well (for example the 2024 ARC-AGI submission by Jack Cole), but still underperform humans. Hence there may be more to the picture [^3] From the perspective of this paper, Active Inference/FEP is primarily a theory of how an agent updates beliefs and selects actions (e.g. via variational free energy minimisation), whereas this paper is more about where and at what level of abstraction adaptation is realised in a system’s stack. Hence I treat the approaches as potentially complementary. An extended comparison and related work is provided in ESM Note S6..
Others seem to suggest biological systems are not more intelligent [^4] In the sense of sample and energy efficiency of adaptation., and the difference is primarily one of inductive bias [chollet2019]. This inductive bias is hard-coded foreknowledge of the environment, which makes it easier to learn and adapt to that environment. To whatever extent this is true, it substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
To reiterate, I want to know why biological self-organisation seems more adaptive [^5] Sample and energy efficient. than contemporary AI. Is there something else biological systems do that allows them to be more ‘intelligent’? Could there be more to the picture than active inference and an inductive bias?
I suggest there is. Computers are organised into layers of abstraction, each layer being the behaviour of the layer below. For example, software runs on hardware. If software is making decisions, then it controls the hardware top-down. In dominant contemporary AI deployments, adaptation is concentrated at a comparatively high level of abstraction (e.g. model parameters and software-level state), while the lower computational stack (instruction set, microarchitecture, hardware) is treated as fixed during deployment. [^6] This is not a claim about a fundamental limitation of computational systems: reconfigurable or self-modifying substrates (e.g. FPGAs, adaptive firmware, agentic materials, or bio-hybrid substrates) can in principle shift adaptation downward. The claim here is about the dominant architectures and workflows in contemporary AI. In contrast, biological self-organisation uses the multiscale competency architecture (MCA). Adaptation takes place at every scale, from cells to organs to organisms [fields2020]. It is ‘scale-free’. The same dynamics are observed at different scales. Control is delegated, in the sense that individual cells pursue goals [^7] As opposed to in a computer where the parts are usually designed to passively await instruction.. Cells are a “soft” material that has agency [man2019, ball2023]. A cell has a perceptual field. A cell can co-operate with other cells in this perceptual field. This allows them to form multicellular networks that have a larger perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Each cell has its own goals and is intelligent in the sense of being able to reach those goals by various means [james1890, ciaunica2023]. However each cell can also be part of a collective of cells that in turn pursues its own collective goals. Smaller collectives can form larger collectives, and the pattern repeats at larger scales. Collectives are nested [ciaunica2021, ciaunica2023b]. Each “thinks” in the enactive sense, interacting with its environment [thompson2007, varela2016, rolla2021]. Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [reid2011].
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent [^8] In the specific efficiency-based sense defined above. than contemporary, predominantly fixed-stack implementations of AI [^9] This is not a claim about all possible future AI. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack.. This is because biological self-organisation delegates adaptation to lower levels of abstraction [^10] For readers who want the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2.. Formally, previous work showed adaptability in embodied systems can be measured by the ‘weakness’ of the constraints they embody [bennett2025d, bennett2024a]. Building on this, I prove a theorem (The Law of the Stack) that the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels. This result shows how we can build more robust and efficient hybrid and cyberphysical systems, laying a foundation for future empirical research. Its applications extend beyond artificial intelligence, for example to human organisational and economic systems.
2 Abstraction, Distribution and Delegation
Conventional computing systems are organised into stacks of abstraction layers. More generally, I will characterise systems along three interacting dimensions: abstraction (how layers relate), distribution (how work is spread within a layer), and delegation (the level of abstraction at which adaptation or control resides). How efficiently a system adapts depends on the interaction between these dimensions.
In proposing Stack Theory [^11] Initially proposed in my PhD thesis [bennett2025thesis]. [bennett2025thesis], I merely extend this idea to systems in general. I’ll begin illustrating what I mean by abstraction, delegation and distribution; dimensions along which one may plot any system. For intuition, I will give examples of limit cases for computational, biological, human organisational and economic systems.
2.1 Abstraction
In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA [^12] This is the software library that allows for the large scale utilisation of GPUs (a type of hardware originally developed for games), for machine learning and artificial intelligence., which runs on a GPU [^13] A graphical processing unit (GPU) is a type of hardware that allows for massively parallel vector processing. It was developed for the purpose of graphics processing in videogames, and later repurposed for machine learning..
GPUGPUCUDACUDAlibrarylibrarymodelmodel
These are all examples of abstraction layers. A programming language is a sort of abstraction layer. For example Python is an abstraction layer over C, in the sense that it is interpreted by a program written in the less abstract C programming language. C in turn depends on a program written in Assembly language, which runs on machine code which is interpreted by hardware. Each layer of language is written in the language of the layer below. This structure allows programmers to work more efficiently, within the confines of a standardised and simplified representation of the system. However, these abstraction layers are incomplete, simplified pictures of the system. They “abstract away” unnecessary details. Most importantly, they are embodied in the layer below. Each language is a subset of the behaviour that the layer below can exhibit. For example, a subset of the possible behaviour a given computer can exhibit would correspond to running Python code, and a broader subset of that computer’s behaviour would necessarily correspond to running C code. Hence each layer in this stack is a subset of the behaviours of the layers below [^14] A higher-level abstraction is not uniquely determined by its lower-level implementation: the same abstract interface can be realised by multiple implementations. The claim here is that higher-level behaviour must be realised by lower-level behaviour, and that different implementations can differ in resource cost (time, energy, information), which matters for adaptability.. This reductionist framing does not deny that useful abstractions can be robust to lower-level variation; invariants can be treated as higher-level task constraints. Even the physical hardware itself can be framed as an abstraction layer over the environment, being a subset of the behaviour of that environment in which the hardware exists [bennett2025thesis].
Think of a stack of abstraction layers as a directed graph. Highly abstract behaviours are a consequence of ‘weaker’, more generalisable goal directed behaviour at lower levels of abstraction. Framed this way, everything is a stack of abstraction layers [^15] This is not an analogical or metaphorical application of computational frames to others systems, as in Bratton’s re-framing of geopolitics [bratton2016]. Rather, this is a more general claim about how all systems work.. For example a collective of cells can enact an organ, and a system of organs can enact an organism [levin2024] which enacts even more abstract behaviours like language or tools. Each of these is an ‘abstraction layer’ in a biological stack [bennett2024a].
cellscellsorgansorgansorganismorganism
2.2 Distribution
In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [delgado1997], which is to say the work is distributed among parts at a level of abstraction. Because of this, looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale. Conversely, rather than looking at low levels of abstraction like cells I can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives at higher levels of abstraction. For example a population of soldiers can be a squad, and if squads co-operate they can be a platoon. Each layer above is the behaviour of a layer below, and the parts get smaller when I go down.
soldierssoldierssquadssquadsplatoonplatoon
In each layer work is distributed among members of the population. If level of abstraction is one dimension along which a system may be plotted, then distribution of work is another. In a computer I can distribute work by splitting it up into sub-tasks and allocating it to many computers.
However, unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective. Looking at a lower level of abstraction is not like ‘zooming in’. I can run software on a solitary computer, or a cluster of computers. For example, Python code is run by an interpreter, which is a C program that in turn is built on Assembly, which is built on machine code that has a fixed interpretation physically hard-wired in the instruction set architecture of the CPU. The scale of the smallest part at each layer of abstraction is the same. Think of one Python program running on one interpreter, on one computer. This illustrates how level of abstraction can vary without affecting the scale or distribution.
2.3 Delegated Control
The aforementioned stack on which Python relies is not a collective. It is not made up of parts capable of functioning independently. They do not have agency. This means control of the system resides at another level of abstraction. If I am programming a computer, then control resides at the level of abstraction at which I am programming. If I am programming in Python, then adaptation occurs at the Python level of abstraction [^16] This is a statement about where learning/adaptation is typically implemented in conventional software stacks. Lower layers can of course carry mutable state (data, caches, firmware state), but are usually not designed to be autonomously goal-directed in the way MCA components are..
Running a program in a distributed system does not always affect this. Distributing work is not the same thing as delegating control or agency. Work can be distributed among many GPUs, but if I remove a GPU from the computer it will not then pursue goals independently. It will sit, inert until once again integrated into a larger system. It requires no incentives. Control is imposed top-down, from higher to lower levels of abstraction. The parts at lower levels have no agency.
Conversely in a biological system each cell is independently capable of goal directed behaviour [ball2023], assuming it remains alive. Control in a complex biological system begins bottom-up, before top-down. Control then propagates in both directions [^17] In the sense that members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with.. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs. Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
A good example to understand the limit cases for delegation is an economy. In an absolutely centrally planned economy, control is concentrated at the highest level of abstraction. One might think of the people in the economy as parts at lower level of abstraction. They do work but have no control. Those people could adapt, but central planner’s edicts constrain them. The opposite extreme would be to have control delegated entirely to the lowest level of abstraction, in anarchy unconstrained by any laws.
The level of abstraction is often related to the granularity at which the system is represented. For example if the level of abstraction at which I view an economic system is humans, then I don’t worry about how work is distributed or control is delegated within those humans. I just concern myself with what systems can be formed of humans, like economies or armies (see fig 1). Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control. Delegating to a lower level of abstraction often requires distributing work among parts at smaller scales [^18] For example, delegating control to the level of individual humans would mean those humans control their own work potential, but not that of others, or else we’d just be delegating to collectives within the human population.. However as the static abstraction layers in a computer illustrate, a lower level of abstraction is not always more granular, and thus work is not always more distributed (see fig 2).
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).
2.4 Reasons to Suspect Delegation Facilitates Adaptation
I will now give a brief explanation of why I believe delegation of control to lower levels of abstraction is relevant to adaptation.
Throughout this subsection I use examples (e.g. distributed drone swarms, mission command, and economies) as intuition-building thought experiments. They are not intended as comparative empirical evidence for the formal claim, which is established later in Theorem 5.1. All else being equal the lower the level of abstraction at which the system can be controlled, the more it can be specialised. For example a program written in C can be far more efficient than a program written in Python [oden2020]. This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C. By programming in C I simply cut out an abstraction layer. This gives the programmer more leeway to optimise the system, or make mistakes. Likewise another way to make computers more efficient is to use a FPGA to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [nguyen2020].
With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt. For example software is nothing more than a state of hardware [^19] That software is a state of hardware has been addressed at length elsewhere [bennett2024a, bennett2025thesis], beyond the scope of this paper.. It is a label I ascribe to the behaviour of hardware. It is a useful simplification that lets us consider what I want hardware to do in the abstract, but working at such a high level of abstraction can cause problems. Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware. The software doesn’t interact with the environment, the hardware does. This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [bennett2024b]. Consequently an ‘objective’ measure of artificial intelligence [leike2015, bennett2024a, bennett2025b] must be ‘enactive’ [thompson2007] in that it hinges not just software but hardware and the local environment as well. To understand the difference between biological systems and artificial intelligence, I must account for the level of abstraction at which an adaptive process takes place.
In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [kuniavsky2010]. For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. It is a single point of failure. In contrast if I can get the swarm to co-ordinate through entirely decentralised control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled [^20] This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [cavagna2010].. In other words, the swarm ‘adapts’ to adverse conditions.
The advantages of delegating control are visible in human organisations as well. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010]. Conceptually, each layer of management is an abstraction layer: when control is delegated, feedback loops shorten because fewer layers sit between the controlling entity and local conditions. This supports the broader hypothesis that delegation is a key ingredient of biological self-organisation that is largely absent from mainstream contemporary artificial intelligence. See ESM Note S7 for a concrete historical illustration (Nordbat 2).
3 What is Adaptation?
I will discuss what it really means to learn and adapt, in terms of cause and effect.
causeeffect
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
not the causeeffectcause
By actively intervening in the environment, a system can experiment to establish cause [^21] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables. This needs only attraction and repulsion from physical states [bennett2023c, bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence. I take this latter approach.
Learning nodes to fit a graph is like figuring out which computer program [^22] Computation in this paper is about constraints, like how slime mold can be made to “compute” by imposing the right constraints on them[nakagaki2000]. I’m not assuming Turing machines. was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^23] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [sole2022]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^24] To the extent that such possible worlds can be differentiated from one another within the subjective confines of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [^25] Formal definitions and proofs for weakness and w-maxing are in [bennett2024appendices, bennett2025thesis]; an extended intuition is in ESM Supplementary Note S3. of the two [bennett2023c, bennett2025d, bennett2025e]. The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b]. This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler policies [^26] Basically minimum description length policies [rissanen1978]., to conform to the terminology used in previous work [bennett2025b, bennett2025thesis] [^27] For formal definitions of learning and efficiency used in this framework, see [bennett2024appendices] and the thesis [bennett2025thesis].. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world [^28] For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture.. Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool [^29] To be exact, a weaker policy completes a superset of tasks, rather than just more tasks. This is because policies are compared in the context of an ostensive definition.. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others [^30] See ESM Supplementary Note S3 for extended discussion of Bennett’s Razor, weakness, and related connections (e.g. maximum entropy).. To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [bennett2022a, bennett2023b]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [bennett2024a]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.
4 Formal Framework and Definitions
The formal definitions below are required for the core claim and for Theorem 5.1. The surrounding narrative provides intuition, but the theorem and its assumptions depend on these definitions. However this paper is based on a formal mathematical framework and it does contain a very simple proof supporting the claim. The formal details are also useful to connect this work here to related mathematical and experimental work [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
Stack Theory describes an adaptive system or computation by the binary constraints or boundary conditions that define it. It formalises embodiment as a choice of abstraction layer: a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis], which are used to express “statements” constraining the body. Policies and tasks are then defined using statements in this embodied language. Delegation, distribution and abstraction are captured by the task hierarchy. A longer conceptual motivation (including the environment/states/programs framing and links to enactivist and pancomputational interpretations) is provided in the Electronic Supplementary Materials (ESM) Note S5 and is not required to follow the theorem below.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A system [ciaunica2023b] that conformed to the free energy principle (FEP) would be a $\mathfrak{v}$-task for which $\langle I,O\rangle$ minimised surprisal. In other words, a $\mathfrak{v}$-task can describe an organism as behaviour that serves homeostatic and reproductive goals.
For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
  1. 1.
    A set $\Phi$ is assumed, whose elements are called states. A declarative program is a set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$. $P=2^{\Phi}$ is the set of all programs. Embodiment is formalised as an abstraction layer, for which I single out a finite set $\mathfrak{v}\subseteq P$ called a vocabulary. A choice of vocabulary $\mathfrak{v}$ implies a formal language
    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}.$
  2. 2.
    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcap l$. A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t. $l_{1}\subseteq l_{2}$.
    The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in L_{\mathfrak{v}}$ has
    $E_{x}=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
    The extension of a set $X\subseteq L_{\mathfrak{v}}$ of statements is
    $E_{X}=\bigcup\limits_{x\in X}E_{x}.$
    The weakness of a policy $x$ is the cardinality of its extension $\lvert E_{x}\rvert$.
  3. 3.
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the set of outputs available given $I_{\alpha}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs. All tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$. A parent is higher level than its children. A parent task is delegated to child tasks, and as each task is embodied this describes control being delegated within a system.
  4. 4.
    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff
    $\pi\in\Pi_{\alpha}=\{\pi\in L_{\mathfrak{v}}:E_{I_{\alpha}}\cap E_{\pi}=O_{\alpha}\}.$
    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}\cap E_{\pi}$. A task is complete if $o\in O_{\alpha}$. Given a task definition $\alpha$ s.t. $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
  5. 5.
    If $\alpha_{1}\sqsubset\alpha_{2}\sqsubset\omega$ then $\alpha_{1}$ contains fewer examples of $\omega$ than $\alpha_{2}$. System $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if it learns $\omega$ from fewer examples. For example $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if $\mathfrak{a}$ learns $\omega$ from either $\alpha_{1}$ or $\alpha_{2}$, while $\mathfrak{b}$ learns $\omega$ from $\alpha_{2}$ but not from $\alpha_{1}$.
  6. 6.
    A causal identity for $\mathfrak{b}$ is a policy classifying the interventions undertaken by $\mathfrak{b}$.
  7. 7.
    $\Gamma_{P}$ contains every task in every vocabulary. For every $\rho\in\Gamma_{P}$ there exists an uninstantiated-task function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}.$ $\lambda_{\rho}$ takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns the highest-level child of $\rho$ that is a $\mathfrak{v}^{\prime}$-task.
  8. 8.
[…]
I set out to compare computers with biological systems. Since both are stacks of abstraction layers, I define multilayer architecture (MLA) below to describe them in comparable terms.
  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ s.t. $\lambda^{i+1}\sqsubset\lambda^{i}$. Lower in the stack is a lower level of abstraction.
  • $\mathfrak{f}:2^{P}\times L_{\mathfrak{v}}\rightarrow 2^{P}$ is a function that takes a vocabulary $\mathfrak{v}$ and a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary
    $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}.$
  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that
    $\mathfrak{v}^{i+1}=\mathfrak{f}(\mathfrak{v}^{i},\pi^{i})\quad\text{and}\quad\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}.$
I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained.
5 Delegated Control and Collective Identity
The vocabulary $\mathfrak{v}$ of an abstraction layer can describe a distributed system. For example, it could describe an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$. The behaviour of two cells as individuals is formalised as two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
$\alpha_{1}$$\alpha_{2}$$\omega$
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
Mechanistically, “negotiation” between levels is represented as constraint compatibility between policy sets. Individual components select policies that remain correct for their local tasks while being consistent with the collective task, and the collective task in turn constrains components by shaping the inputs they experience. When the intersection of viable policies becomes empty, no collective policy exists and collective identity collapses (a failure mode analysed in Section 6). In hybrids, the designer mediates this negotiation by adjusting boundary conditions (which reshape the component policy sets) rather than by directly specifying a single low-level policy.
5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [bennett2024a]. Put another way, ‘organ’ is a label I ascribe to the the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone; ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
The cells are tasks. For tasks at higher levels of abstraction, capacity for w-maxing depends on lower levels. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ [^31] I can substitute $o$ for a declarative program $f$ that has the same extension, because that means $f$ and $o$ are true of the same states, having the same ‘truth conditions’ so to speak.. Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. I can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define ‘correctness’ at every level. At each successive level, ‘correct’ gets more specific. Because everything that happens in a higher level is in some sense correct according to lower levels, each higher level must more tightly constrain what is considered to be correct behaviour if that higher level is to exist in a meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In purely narrative terms, the implication of all this is that adaptation at higher levels depends on adaptation at lower levels. Below I state The Law of the Stack with a brief proof sketch; a full proof is provided in the Electronic Supplementary Materials.
Theorem 5.1 (The Law of the Stack).
Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$. Fix any $i<n$, and write
$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i}),\qquad\gamma_{i+1}=\lambda_{i+1}(v_{i+1}).$
Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense:
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
Equivalently,
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|E_{\pi_{i}}|$, i.e. weakening the policy at the lower layer.
Proof sketch.
Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $E_{\pi_{i}}$ under the map $o\mapsto\bigcap o$, hence $|v_{i+1}|\leq|E_{\pi_{i}}|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|}$. ∎
Intervening in the stack limits the extent to which control can be delegated, and thus limits capacity for w-maxing or MCL. It makes lower layers of abstraction static in the sense that they no longer change to facilitate adaptation at higher levels. By Theorem 5.1, any increase in achievable utility at layer i+1i{+}1 requires (at least) an increase in the extension size of the realised policy at layer $i$, i.e. support for weaker constraints below.
5.2 Human Abstractions
Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent behaviours of that MCA.
cellscellshumanshumansabstractionsabstractions
When we humans create a computer, we create a machine that embodies a subset of our more abstract behaviours. An engineer hard-wires those behaviours into the instruction set architecture (ISA), and we all then ascribe meaning to that behaviour. Two results can look the same at the higher level (e.g. the number two, spoken by a human or displayed on a screen), but be fundamentally different at the lower level (e.g. resources consumed, data format etc). The results are only the same if one ignores the lower levels of abstraction.
Software might be agentic, but computer hardware is passive. Control is top-down. Once I have built an ISA, it exists in the environment independently of the human MCA from which it emerged. It is static. It cannot adapt, which is why there is demand for field programmable gate array (FPGA) [^32] An FPGA is like a CPU that can be reconfigured for specialised applications..
humanshumansISAISAabstractionsabstractions
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [levin2021, davies2011]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
In the context of my formalism, a state analogous to cancer occurs when a system is ‘over-constrained’ so that no collective policy exists, while the capacity for independent action remains. For example, assume a collective represented by a task. Adverse environmental conditions are represented as a task for which fewer correct policies exist than less adverse environmental conditions. Likewise excessive top down control can limit adaptability, and thus the number of correct policies which exist. If the constraints are too tight, then no correct policies will exist and the collective’s identity will be lost.
The set of correct policies shrinks as task difficulty increases, meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. If the policy at a lower level i1i-1 changes such that the task at level $i$ becomes ‘harder’333333i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in there being no correct policy at level $i$. If there are no longer any correct policies at level $i$, then the only way the system can continue to function is for part of the collective to ‘splinter’ or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind. Say I have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(\mathfrak{v}^{i-1}_{\mathfrak{a}},{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to ‘selective forgetting’ [bennett2023c] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, I would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that I obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of ‘correct’ policies to choose from.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. To summarise; a state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
7 Hybrid Agency as Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input–output relation at the interface.
The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input–output mapping at the interface.
According to The Law of the Stack, achieving higher utility at an interface layer requires sufficiently weak policies at lower layers. Therefore, boundary conditions should be as weak (least exclusionary) as possible while still making the target behaviour achievable and preserving a non-empty intersection of component policy sets (collective identity). Overly tight constraints can collapse that intersection and force failure, while overly loose constraints may fail to stabilise any desired collective policy. A schematic formalisation of this boundary-condition design view is provided in the ESM.
8 Conclusions and Future Research
In this paper I used Stack Theory to describe MCAs as stacks of abstraction layers. In this view policies are embodied at each layer, approximating causality in service of adaptation, and cause is best approximated by weaker policies (w-maxing) [bennett2023c]. Formally, I prove a theorem (The Law of the Stack 5.1) showing w-maxing in higher layers requires sufficient w-maxing in lower layers. One reason artificial intelligence tends to be less adaptable is because it rests upon an inflexible stack of human abstractions. This claim is directed at contemporary, predominantly fixed-stack AI architectures. It is not a fundamental limitation on future computational or hybrid systems.
My claim is that, to build more robust systems, control should be delegated to the greatest extent possible. Obviously this must be balanced against desired functionality. Policies must be aligned. The system must still discriminate between correct and not. In the earlier sections I discussed the doctrine of mission command. What I am suggesting is that the concept of mission command should be applied to systems in general. At least, to the extent that adaptability is desired. Put another way, any constraints which are not strictly necessary to maintain the integrity of the system should be eliminated [^34] What counts as “strictly necessary” is relative to design goals and value judgements; see ESM Supplementary Note S4.. Humans tend to prefer ‘additive’ solutions [adams2021], by which I mean if a problem can be solved by adding a rule or subtracting a rule, we humans tend to add a rule. To effectively delegate control, we should subtract rules instead.
There are also software-first approaches that approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6).
Future research might consider how adaptive hardware can be designed, to delegate control. Multilayer-causal-learning (MCL) provides a formal basis for ‘soft robotics’ [man2019]. Soft robotics is a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, providing a formal foundation for soft robotics. Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction. They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, what might be required is a sort of artificial life, with small parts interacting form complex systems from the small scales and low levels of abstraction up [suzuki2006, ikegami2008, ikegami2007simulating].
As for the inductive bias mentioned in the introduction [chollet2019], it boils down to the choice of abstraction layer [bennett2024a]. I have described a means of learning a suitable abstraction layer. In theory, a multilayer-causal-learning system should be sufficient to solve ARC-AGI benchmark problems in terms of both sample and energy efficiency. Future research should test this empirically through (i) agent-based simulations of collectives with tunable delegation and abrupt environment shifts, measuring sample/energy efficiency and robustness; (ii) comparative benchmarks in which adaptation is confined to high-level software state versus systems with reconfigurable substrates (e.g. adaptive firmware or FPGAs); and (iii) hybrid experiments in which boundary conditions on biological components are systematically varied to quantify trade-offs. A more detailed set of candidate evaluation designs is provided in ESM Note S8.
Beyond hardware, cyberphysical systems in general can be evaluated in terms of abstraction, distribution and delegation. I might consider how states analogous to cancer could arise in human organisational structures, or how control can be delegated throughout an organisation to facilitate robustness and efficiency. For example, this could be applied to facilitate transformational bottom-up rather than incremental top-down adaptation to environmental changes.
Free-market trade is a great example of how delegation can facilitate adaptability in a large collective. As Hayek pointed out [hayek1945use], not all of the information is available to a central planner and so control is best delegated to the individual. The accumulation of resources within my formalism is implicit, rather than explicit. I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e. does it achieve homeostatic and reproductive goals in whatever amount of time care about?), rather than how it evolves over time as a dynamical system. In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there. Future research might formalise the accumulation of resources explicitly. One way to describe an evolving sequence of tasks over time could be something like a stack. To speculate about what my results imply for delegated systems like an economy, under less than ideal conditions resources may become concentrated in the hands of a few. The resulting system then delegates work, but not control. Those parts deprived of control may then not be able to construct a policy that aligns with the collective, and so they may splinter off and become ‘cancerous’. To prevent this, mechanisms can be put in place to preserve decentralisation without the need for intervention from outside the system (e.g. government intervening to break up a monopoly). Each system could be located along axes of redistribution and delegation, like an alternative political compass. Just enough redistribution of control and the system remains delegated. However, too much redistribution could have a similar effect, preventing parts from adapting to local conditions. An example of a mechanism that can effectively preserve delegation of control is adaptive rubber banding [mi2022]; a game mechanic (e.g. in Mario Kart) where success accrues penalties as well as resources, keeping the game competitive.
Another possibility is AI regulation. For example the UK government recently initiated “Safeguarded AI” [dalrymple2024b, dalrymple2024a], a project that aims to construct an AGI [goertzel2014] of level ASL-4 [anthropic2023] in a cyberphysical system, to safeguard against rogue AI. The idea of a “rogue AI” from within a cyberphysical system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective. AI regulations may over-constrain systems and trigger the very cancerous state they seek to prevent. To reiterate the earlier point about subtractive solutions [adams2021], any guardrails which are not strictly necessary must be eliminated to make room for those that are.
In summary, I have described systems in terms of delegation and abstraction. I proved that adaptability at higher levels of abstraction is constrained by adaptability at lower levels (The Law of the Stack). I have argued that one reason biological systems are more ‘intelligent’, meaning sample and energy efficient, is because they delegate adaptation down the stack. This Stack Theoretic perspective lays a foundation for future empirical research, showing how we might build more robust systems of any kind.
9 Electronic Supplementary Materials
9.1 Mathematical Results
The Law of the Stack
Theorem 9.1 (The Law of the Stack).
Assume the formal setting defined in the main text: finite vocabulary $v\subseteq P$, language $L_{v}=\{\,\ell\subseteq v:\bigcap\ell\neq\emptyset\,\}$, extensions $E_{\pi}=\{\,y\in L_{v}:\pi\subseteq y\,\}$, $v$-tasks $\gamma=\langle I_{\gamma},O_{\gamma}\rangle$, correct policies $\Pi_{\gamma}=\{\pi\in L_{v}:E_{I_{\gamma}}\cap E_{\pi}=O_{\gamma}\}$, and utility $\epsilon(\gamma)=\max_{\pi\in\Pi_{\gamma}}(|E_{\pi}|-|O_{\gamma}|)$. Let $\langle\lambda_{0},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
Fix any $i<n$, and define $\gamma_{i+1}=\lambda_{i+1}(v_{i+1})$ where $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$. Then
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
Equivalently,
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
Thus, higher utility at layer i+1i\!+\!1 requires (at least) a weaker policy at layer $i$ in the sense of larger extension cardinality.
Proof.
We proceed by a sequence of elementary bounds that make the layer-to-layer dependency explicit.
Lemma 1 (Languages are bounded by the power set).
For any finite vocabulary $v$, the language $L_{v}$ satisfies $|L_{v}|\leq 2^{|v|}$.
Proof. By definition, $L_{v}\subseteq\mathcal{P}(v)$ because every $\ell\in L_{v}$ is a subset of $v$. Since $|\mathcal{P}(v)|=2^{|v|}$ for finite $v$, we have $|L_{v}|\leq 2^{|v|}$. $\square$
Lemma 2 (Extensions live inside the language).
If $\pi\in L_{v}$, then $E_{\pi}\subseteq L_{v}$, hence $|E_{\pi}|\leq|L_{v}|$.
Proof. By definition $E_{\pi}=\{y\in L_{v}:\pi\subseteq y\}$, so $E_{\pi}\subseteq L_{v}$. Taking cardinalities yields $|E_{\pi}|\leq|L_{v}|$. $\square$
Lemma 3 (Utility implies a lower bound on language size).
Let $\gamma$ be a $v$-task. Then
$\epsilon(\gamma)+|O_{\gamma}|\leq|L_{v}|.$
Proof. Because $v$ is finite, $L_{v}$ is finite, hence $\Pi_{\gamma}\subseteq L_{v}$ is finite, so the maximum defining $\epsilon(\gamma)$ is attained by some $\pi^{\star}\in\Pi_{\gamma}$. By definition,
$\epsilon(\gamma)=|E_{\pi^{\star}}|-|O_{\gamma}|\quad\Rightarrow\quad\epsilon(\gamma)+|O_{\gamma}|=|E_{\pi^{\star}}|.$
By Lemma 2, $|E_{\pi^{\star}}|\leq|L_{v}|$. Combining yields the claim. $\square$
Lemma 4 (The abstractor output vocabulary is an image of an extension).
Let $T:L_{v_{i}}\to P$ be the map $T(o)=\bigcap o$. Under the MLA definition,
$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})=\{\,T(o):o\in E_{\pi_{i}}\,\},$
so $v_{i+1}$ is the image of $E_{\pi_{i}}$ under $T$. Consequently,
$|v_{i+1}|\leq|E_{\pi_{i}}|.$
Proof. This is just re-writing the definition of $\mathfrak{f}$ in functional form, and then applying the general fact that for finite sets $A$ and any function $g$, $|g(A)|\leq|A|$. $\square$
Conclusion.
Apply Lemma 3 to $\gamma_{i+1}$ in vocabulary $v_{i+1}$ to obtain
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|.$
Apply Lemma 1 to bound $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$, giving
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|v_{i+1}|}.$
Finally apply Lemma 4: $|v_{i+1}|\leq|E_{\pi_{i}}|$. Since $2^{x}$ is monotone increasing,
$2^{|v_{i+1}|}\leq 2^{|E_{\pi_{i}}|}.$
Combining the last two displays yields
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|},$
and the logarithmic form follows by applying $\log_{2}$ (also monotone). ∎
Remark. In the thesis version of this result, the intuition is phrased as: larger vocabularies at layer i+1i{+}1 permit weaker (more general) policies, hence more adaptation at higher layers. The proof here makes the dependency explicit by bounding the achievable utility at layer i+1i{+}1 as a monotone function of the extension size $|E_{\pi_{i}}|$ at layer $i$, without requiring any additional monotonicity assumptions about $\epsilon(\lambda(\cdot))$ under set inclusion.
Hybrid case study formalisation (Pong organoid)
Let $C$ be a set of cells. For each cell $i\in C$, let $\Pi_{i}$ denote the set of viable policies for that cell under some interface/environmental conditions. The set of viable collective policies for the organoid is $\Pi_{C}=\bigcap_{i\in C}\Pi_{i}$. A conventional computer interface (electrodes, stimulus generator, game process) imposes boundary conditions $c$ that restrict which policies remain viable; write this restriction schematically as an operator $R_{c}$ such that $\Pi_{i}^{(c)}=R_{c}(\Pi_{i})$ and $\Pi_{C}^{(c)}=\bigcap_{i\in C}\Pi_{i}^{(c)}$. The design objective is to choose $c$ so that (i) $\Pi_{C}^{(c)}\neq\emptyset$ and (ii) the remaining collective policies implement a target output set $O_{\text{pong}}$ at the interface. In this view, building the hybrid is a constraint-design problem: impose the weakest (least exclusionary) boundary conditions that make $O_{\text{pong}}$ achievable while preserving non-empty intersection (collective identity).
9.2 Supplementary Notes and Extended Discussion
S1. Working definition of intelligence
There are many definitions of intelligence, and the merits of each are beyond the scope of this paper. Mine is based on [bennett2025thesis], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation. It is similar to Wang’s definition of intelligence as adaptation with limited resources [wang2019], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance. By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems. This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
S2. Levels of abstraction beyond computing
The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [bennett2025thesis]. A higher level of abstraction is the behaviour of the layers below. For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
S3. Weakness and w-maxing (extended intuition)
Weakness is formally defined in the appendix. Related work proved w-maxing to be optimal regardless of the abstraction layer [bennett2025e, bennett2023b, bennett2024a, bennett2025thesis]. In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness. If $x$ is weaker than $y$, that means it commits to less, but achieves the same end. For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [bennett2023b]). This is a bit like the principle of maximum entropy [jaynes1957a, jaynes1957b], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [bennett2024b].
For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [bennett2024appendices]. It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
S4. Normativity and “strict necessity” in design
As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is. “Strictly necessary” is a value judgement, or an implicit choice of what ought to be. In the thesis [bennett2025thesis] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment. These transitions preserve some aspects of that environment, while destroying others. Here however I am more concerned with the design of systems which serve a human purpose. Strictly necessary is a matter of desired functionality.
S5. Conceptual motivation for the formal setting
Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [bennett2025thesis]. The full justification for these definitions is out of scope and addressed at length elsewhere [bennett2024a, bennett2025thesis, bennett2025e]. However I will provide a brief overview. To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer. From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite). I call this bottom layer the environment. Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction. Hence there must be a set of states where each state is one point of difference [^35] Nothing is assumed about what these states might ontologically be or epistemologically contain.. Because these states must be mutually exclusive, this frames time as difference. Each possible sequence of states is a timeline or possible world. Truth is defined ontologically, relative to states. A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state. Every aspect of the environment is just a set of programs. Everything that exists is a set of facts, and each timeline preserves different sets of facts. A timeline is like an opinion; as time progresses some aspects of the environment are preserved and others are destroyed. Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [bennett2025thesis]. If the set of states is infinite then there are infinite possible worlds. Assume I have a robotic system. I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary. Everything the abstraction layer can express is a subset of its vocabulary. For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’. Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language. I do not mean like the words humans use to communicate. I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [nakagaki2000] [^36] Using this frame cognition can be formalised in terms that align with both enactivist [thompson2007] and pancomputationalist [piccinini2021] perspectives. Hence, cognition within Stack Theory is also called Pancomputational Enactivism[bennett2024a].
S6. Extended related work on delegated software architectures and Active Inference/FEP
There are examples of systems that take something like the present approach, albeit within the confines of software abstraction layers.
Federated machine learning delegates control to some extent (e.g. shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [nori2025federated]. Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [thorisson2012, goertzel2014]. The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [nivel2013] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [thorisson2018cumlearncause], “growing” knowledge in a collective meta-model through cumulative continuous learning [thorisson2019cumlearning], and explicitly targeting adaptability [thorisson2014autonomousNatCom, thorisson2020, eberding2024, sheikhlar2024] [^37] See also http://www.openaera.org.. ActPC-Chem [goertzel2024] is a recent theory behind the distributed Hyperon architecture [goertzel2023] that was influenced by theoretical and methodological insights from AERA.
Active Inference/FEP in relation to MCL.
Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [Friston2010, friston2013, parr2024]. The present paper addresses a different axis: it states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1). In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer. An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers; conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
Potential complementarity and open questions.
Several lines of work suggest points of contact, without implying formal equivalence. ActPC-Chem frames delegation in explicitly Active Inference terms [goertzel2024, Friston2010, friston2013]. “Scale-free” active inference proposes that similar dynamical principles repeat across scales [friston2023b, friston2024a], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity. Delegated [ororbia2022, ororbia2023, ororbia2023b] or federated [friston2024b] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale. Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g. the Landauer bound) by reducing memory–processor separation and associated bottlenecks [ororbia2024mortalcomputationfoundationbiomimetic, aurell2012, zou2021], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead. However, the theories still differ in their primitives (e.g. tasks/weakness versus generative models/free energy), and it remains an open question whether weakness (w-maxing) can be recovered as, or justified by, a principle within the Active Inference apparatus. Establishing such a mapping is an important direction for future work.
S7. Mission command example (Nordbat 2)
A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [ingesson2016].
S8. From analogy to evidence: candidate empirical and simulation tests
The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence. To turn the delegation hypothesis into a research programme, one can state testable predictions and compare systems under controlled variation of delegation depth.
1. Agent-based simulations with tunable delegation.
Construct an agent-based model of a collective where (i) individuals have local perception and local policy spaces, (ii) a collective objective is defined at a higher level, and (iii) a “delegation parameter” controls how much decision-making is centralised versus left to individuals (e.g. by constraining local policy spaces or routing decisions through a leader). Expose the collective to non-stationary environments (distribution shifts, component failures) and measure: sample efficiency (data required to regain acceptable performance), robustness (graceful degradation under failures), and identity preservation (whether the collective policy remains coherent rather than splintering into incompatible sub-collectives). Theorem 5.1 predicts that as delegation is pushed upward (lower layers become static or over-constrained), achievable utility at higher layers should become bounded by the reduced weakness of the lower-layer realised policies.
2. Comparative benchmarking across stack depth.
Compare learning systems that differ only in where adaptation is permitted. For example, keep a fixed high-level learning rule but vary whether lower layers are static (fixed hardware / fixed microcode) versus adaptive (e.g. reconfigurable hardware, adaptive firmware, or reconfigurable substrates). Measure sample and energy efficiency on tasks with distribution shift and compositional generalisation (e.g. ARC-style tasks). This isolates the “where in the stack” hypothesis from the choice of algorithm at the top.
3. Hybrid boundary-condition sweeps.
For hybrid agents such as simulated systems of human and AI agents, where the goal is to “align” the AI agents with goal goals, systematically vary the boundary conditions (stimulation regimes, feedback mappings, resource constraints) and quantify the trade-off between task performance and preservation of collective identity (non-empty intersections of viable policy sets, as in the organoid formalisation above). The design expectation is that weaker (less exclusionary) boundary conditions should preserve a larger viable policy space while still enabling the desired interface-level behaviour.
4. Interaction with Active Inference/FEP.
Implement Active Inference-style control at one or more layers, but run the same comparison as above: (i) active inference on a fixed substrate versus (ii) active inference on a substrate that permits deeper delegation (adaptation in lower layers). This helps disentangle improvements due to the inferential objective from improvements due to stack-level delegation.
These designs are intended as starting points rather than definitive protocols; the unifying aim is to replace analogy with controlled variation of delegation, and to measure adaptability in the efficiency-of-adaptation sense used throughout this paper.